4.+-2 points Tipler6 28.P.045.soln A 2.00 cm by 1.50 cm rectangular coil has 380 turns and rotates in a region that has magnetic field of 0.360 T (a) what is the maximum emf generated when the coll rotates at 60 rev/s? (b) What must its angular speed be to generate a maximum emf of 118 V? revis eBook
Pelase HELP!
Answer #1

e = NBAw

N = 380 turns

A = 2 x 1.5 x 10^-4 m^2

B = 0.360 T

w = 2pi*f = 2 * pi * 60

e = 15.47 V

part b )

e = NABw  

w = e/NAB

w = 118/(380 * 2 * 1.5 * 10^-4 m^2 * 0.360)

w = 2875.24 rev/s